OpenStudy (anonymous):

Algebra: The price of an item is increased by 10% and then by another 10% and is now $242. What was the original price? (algebraic solution)

OpenStudy (anonymous):

Well we have the final price at 242. What was its price BEFORE the second markup? Premarkup price = 242 / 1.1 Then we take that price, perform the same procedure on it, and BAM! Your answer.

OpenStudy (saifoo.khan):

Ops. i read the opposite. D:

OpenStudy (anonymous):

successive increase/decrease question.

OpenStudy (lgbasallote):

let's say original price is x so increase by 10% will be x + 0.10x simplifying is 1.10x then you get the 10% of that so \[\large 1.10 x + (1.10x \times 0.10x) = 242\]

OpenStudy (lgbasallote):

lol now that is how it should be...

OpenStudy (anonymous):